MOA-II Obsevation Log Monthly Summary 2009/04

Return to index

Previous month Next month

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB17619-179883863231196080
LMC3886-398096453139460
SMC818-8269186000
ALERT-000
GRB63-7917171560
TOTAL---5083719343100 (95h00m1100s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
2211113383333348562083243204552
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
12731051511019945454480


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
151515141517967873271819
Field13Field14Field15Field16
242700


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
990000000000


SPECIAL TARGET NAME

CategoryTargetNrun
GRBGRBunknowndes117


CCD TEMPERATURE Previous month Next month




VACUUM Previous month Next month




TELESCOPE TEMPERATURE Previous month Next month




COMMENTS
2009/04/1: there was a GRB observed tonight. / night time = 38100(sec)

2009/04/2: There are busses of people comming for the 100 hours of astronomy / night time = 38280(sec)

2009/04/3: live Webcast tonight . / night time = 38460(sec)

2009/04/4: / night time = 38640(sec)

2009/04/5: / night time = 38880(sec)

2009/04/6: / night time = 39000(sec)

2009/04/7: no observation / night time = 39720(sec)

2009/04/8: Some holes but not for long enough / night time = 39360(sec)

2009/04/9: / night time = 39540(sec)

2009/04/10: Full moon. follow-up lmc7-R-10-4396 / night time = 39720(sec)

2009/04/11: / night time = 39900(sec)

2009/04/12: 300sec DARK have a pattern because of CCD is too hot (T=-70) / night time = 40080(sec)

2009/04/13: non / night time = 40200(sec)

2009/04/14: moon was close to GB / night time = 40380(sec)

2009/04/15: No observation / night time = 40560(sec)

2009/04/16: non / night time = 40680(sec)

2009/04/17: none / night time = 40860(sec)

2009/04/18: no obs. / night time = 41040(sec)

2009/04/19: Paul's blue LED was on when DARK was taken. / night time = 41220(sec)

2009/04/20: / night time = 41340(sec)

2009/04/21: seeing was 1.5arcsec / night time = 41520(sec)

2009/04/22: V band observations / night time = 41640(sec)

2009/04/23: none / night time = 41820(sec)

2009/04/24: none / night time = 41940(sec)

2009/04/25: no obs. / night time = 42120(sec)

2009/04/26: We changed vacuum sensor. / night time = 42240(sec)

2009/04/27: We took Flat images (F2764-F2799, F2802). Be aware not to use F2787 which has low-gain image of chip7. / night time = 42360(sec)

2009/04/28: None. / night time = 42540(sec)

2009/04/29: No obs. / night time = 42660(sec)

2009/04/30: None. / night time = 42780(sec)

TOTAL NIGHT TIME: 1217580(sec)

OPERATING RATE: 40.4 %
GB: 26.7 %
LMC: 12.9 %
SMC: 0.6 %
ALERT: 0.0 %
GRB: 0.2 %

Return to index

Previous month Next month